E19.2 (LO 1,2) (Two Differences, No Beginning Deferred Taxes, Tracked through 2 Years) following information is available for Wenger Corporation for 2019 (its first year of operations). The 1. Excess of tax depreciation over book depreciation, $40,000. This $40,000 difference will reverse equally over the years 2020-2023 2. Deferral, for book purposes, of $20,000 of rent received in advance. The rent will be recognized in 2020. 3. Pretax financial income, $300,000. 4. Tax rate for all years, 20%. Instructions a. Compute taxable income for 2019. b. Prepare the journal entry to record income tax expense, deferred income taxes, and income taxes payable for 2019. c. Prepare the journal entry to record income tax expense, deferred income taxes, and income taxes payable for 2020, assuming taxable income of $325,000.
Answer #1
1
Pretax financial income 300000
Less: Excess Depreciation (40000)
Add: Rent received in advance 20000
Taxable income 280000
2
Account Titles and Explanation Debit Credit
Income tax expense 60000
Deferred tax asset 4000 =20000*20%
        Deferred tax liability 8000 =40000*20%
        Income taxes payable 56000 =280000*20%
3
Account Titles and Explanation Debit Credit
Income tax expense 67000
Deferred tax liability 2000 =8000/4
        Deferred tax asset 4000
        Income taxes payable 65000 =325000*20%